Hello math friends:)I cannot set this one up either...The product of two consecutive integers is three less than three times their sum...is it
x(x+1)=3(x+x+1) -3? I tried a lot of ways and cannot get 2 consecutive numbers that work
:(

Answer by Theo(13342)

your formula looks good.
i did the following:
x*(x+1) = 3*(x+(x+1))-3
simplify to get:
x^2 + x = 3*(2x+1)-3
simplify further to get:
x^2 + x = 6x + 3 - 3 which becomes:
x^2 + x = 6x
if you divide both sides by x, you get x = 5
if you subtract 6x from both sides of the equation, you get:
x^2 - 5x = 0 which you factor to get:
x * (x-5) = 0 which results in:
x = 0 or x = 5
0 doesn't work, but 5 does.
5 is the number.
6 is the next consecutive number.
5*6 = 30
3 * (5 + 6) - 3 = 3 * 11 - 3 = 33 - 3 = 30
answer checks out so we're good.
answer is 5 and 6.